Invest $30,000 in a special package of bonds, stocks, and real estate and we guarantee you will earn $4000 this year. The stocks have an expected growth of 30%, the bonds 25% and real estate 10%. How much money can you invest in each type of investment to receive the $4000?
Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)

the stocks 30%, the bonds 25% and real estate 10%.
6x + 5x + 2x = $4000
x = $307.69
307.69 * 6 = $1846.14 stocks
307.69 * 5 = $1538.45 bonds
307.69 * 2 = $615.36 real estate
Total: $3999.97 0r $4000 rounded

Answer by ikleyn(52852)

.
Invest $30,000 in a special package of bonds, stocks, and real estate and we guarantee you will earn $4000 this year.
The stocks have an expected growth of 30%, the bonds 25% and real estate 10%.
How much money can you invest in each type of investment to receive the $4000?
~~~~~~~~~~~


It is a strange  " problem ",  more like an advertisement,  mistakenly submitted to this forum.

If considered for  3  unknowns,  it provides only two equations for three unknown quantities,
leaving possibility for infinitely many solutions.


So,  I am not inclined to consider it seriously.


//////////////////


Comment from student : Umm it's a real problem that you are incorrect.
It's a matrix problem and all you gave x answers and not x, y, z answers.
This makes me think you're not a qualified math instructor.



My response :   you may think everything you want.
                         Of it,  your problem does not become better.


Do you see, when you post the problem, posed incorrectly
and do not understand that you are wrong,
your opinion is not interesting to anybody else at this forum, except yourself.


To make right and solid suggestions, much higher level of knowledge is required,
than what you demonstrate in your post.